PROBLEM 1. The following data came from Pio Corporations shareholders equity: Share premium-Treasury Shares Transactions P
Answer #1

From the data given of shareholders equity,

It is implied that Pio Corporation currently holds 1000 treasury shares, means 2000 shares out of originally held 3000 shares has been sold. Moreover, premium is also reflecting for 2000 shares.

Therefore,

cost per treasury share =20000/1000 =$20

premium per treasury share =6000/2000 =$3

Total selling price per share =20+3 =$23